OpenStudy (anonymous):

A shipping container is in the shape of a rectangular box that has a length of 10x+15 units, a width of 5x units, and a height of 2 units write an expression that can be used to find the volume of the box and then factor the expression completely.

OpenStudy (anonymous):

I think the first part of the answer is 5(2x+1x+3)*2 am i right?

OpenStudy (btaylor):

V = (L)(W)(H) V = (10x+15)(5x)(2) V = 100x^2 +150x V = 50x(2x+3)

OpenStudy (anonymous):

How does (5x)(2)=150x

OpenStudy (anonymous):

@BTaylor How does (5x)(2)=150x

OpenStudy (btaylor):

it doesn't. I skipped a step: (10x + 15)(5x)(2) (10x + 15)(10x) 100x^2 +150x get it now?

OpenStudy (anonymous):

Yes I do but then from there how would I factor it?

OpenStudy (anonymous):

You can factor 10 x

OpenStudy (btaylor):

both terms have an x, and are multiples of 50. So you can pull out a 50x, and you get 2x+3 as the other factor.

OpenStudy (anonymous):

100x^2 +150x=10x*(????)

OpenStudy (anonymous):

@BTaylor what happens to the x's when you factor the equation?

OpenStudy (anonymous):

50x*(2x+3)

OpenStudy (anonymous):

It looks to me like it would be 2x^2+3x??? Tell me if I'm wrong. I'm trying to teach myself algebra.

OpenStudy (btaylor):

there is an x in each term, so divide both terms by x. You get 2x + 3

OpenStudy (anonymous):

so you would divide each term by 50x then?

OpenStudy (btaylor):

yes. That is called factoring out the greatest common factor (GCF)

OpenStudy (anonymous):

Thank you so muc.